Jones and Smith together weigh 340 pounds. Find their weights if they balance a lever when Jones sits 5 feet from its fulcrum on one side and Smith sits 6 feet from the fulcrum on the other side.
: Jones and Smith together weigh 340 pounds. Find their weights if they balance a lever when Jones sits 5 feet from its fulcrum on one side and Smith sits 6 feet from the fulcrum on the other side.

Answer by scott8148(2761)

"Jones and Smith together weigh 340 pounds" __ J+S=340 __ J=340-S

"they balance a lever when Jones sits 5 feet from its fulcrum on one side and Smith sits 6 feet from the fulcrum on the other side"

when they balance, the products of their weights and distances from the fulcrum are equal __ 5J=6S

substituting __ 5(340-S)=6S __ 1700-5S=6S __ 1700=11S __ 1700/11=S

substituting __ J=340-(1700/11) __ J=2040/11